OpenStudy (anonymous):

A carton contains 12 light bulbs of which 3 are defective. if a random sample of 4 light bulbs is chosen without replacement, what is the probability distribution of defective light bulbs?

OpenStudy (anonymous):

Pr(defective bulbs)=3/12=1/4 Pr(light bulbs)=9/12=3/4

OpenStudy (anonymous):

can someone please help me find the answer?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

probability distribution means you want the probability that 0, 1, 2, 3, or 4, bulbs are defective right?

OpenStudy (anonymous):

probability that none are defective is \[\frac{9}{12}\times \frac{8}{11}\times \frac{7}{10}\times \frac{6}{9}\] is one way of computing it

OpenStudy (anonymous):

i didnt get it. what do they mean by the sentence "chosen without replacement"?

OpenStudy (anonymous):

it means pick one put it aside then pick another one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

here is another way to do it

OpenStudy (anonymous):

could u please teach me both ways

OpenStudy (anonymous):

probability distribution means you want the probability that 0, 1, 2, 3, or 4, bulbs are defective right?

OpenStudy (anonymous):

why 4?

OpenStudy (anonymous):

the number of ways you can pick 4 out of 12 is called "12 choose 4" written either as \[\binom{12}{4}\] or as \[_{12}C_4\] and computes as \[\binom{12}{4}=\frac{12\times 11\times 10\times 9}{4\times 3\times 2}\] \[=11\times 5\times 9=495\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh duh, yeah, you are right it can only be 0, 1, 2, or 3, i was being stupid

OpenStudy (anonymous):

ok so \(495\) will be the denominator for all your answers

OpenStudy (anonymous):

now lets compute the probability you pick no defective ones that is 3 out of the 9 not defective, so \(\binom{9}{3}\) is your numerator

OpenStudy (anonymous):

therefore \[P(x=0)=\frac{\binom{9}{3}}{\binom{12}{4}}\]

OpenStudy (anonymous):

could you pless tell me by what they mean by "probability distribution of defective light bulbs"?

OpenStudy (anonymous):

please*

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you pick 4 light bulbs total out of the 12 right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

out of those 4, how many can be defective? you answered that already, it can be 0, 1, 2, or 3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if you put say \(x\) as the number of defective light bulbs chosen then there are 4 possibilities for \(x\) it could be \[x=0,x=1,x=2,x=3\]

OpenStudy (anonymous):

the "probability distribution of defective light bulbs" are the four numbers \[P(x=0), P(x=1), P(x=2), P(x=3)\]

OpenStudy (anonymous):

im writing these down as notes :)

OpenStudy (anonymous):

in other words you have to compute four probabilities the probability none are defective the probability one is defective the probability two are defective the probability three are defective

OpenStudy (anonymous):

man your the best :)

OpenStudy (anonymous):

lol thanks now we still have to compute the 4 numbers, right?

OpenStudy (anonymous):

ill do it myself and you check my answer if thats alright?

OpenStudy (anonymous):

yeah that if fine but there is a quick way to do it, more or less

OpenStudy (anonymous):

so if you are knocking yourself out, you are doing too much work let me know what you get

OpenStudy (anonymous):

great. ill finish writing this down and ill get back to you my friend

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can we use the general formula to get the answer? through k, n, p?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you could if the bulbs were being replaced, because then the probabilities would always be the same but they are NOT being replaced (remember is says "without replacement") so the probabilities change

OpenStudy (anonymous):

by which i mean, if you pick a defective one first, the probability the next one is defective is no longer \(\frac{3}{12}\) it is now \(\frac{2}{11}\)

OpenStudy (anonymous):

i'll be happy to show you the more or less snappy way to do this if you like

OpenStudy (anonymous):

please im finding it a bit confusing

OpenStudy (anonymous):

ok lets pick a middle one, say \[P(x=2)\] i.e. the probability that exactly 2 are defective we keep our denominator as \(_{12}C_4=495\)

OpenStudy (anonymous):

now we are picking 4 so 2 defective and 2 not defective the number of ways to pick the 2 out of 3 defective ones is \(_3C_2=3\) the number of ways to pick the 2 out of 9 not defective ones is \(_9C_2=36\)

OpenStudy (anonymous):

answer is therefore \[P(x=2)=\frac{3\times 36}{495}\]

OpenStudy (anonymous):

we can write the answer immediately although we still have to compute the numbers 2 out of 4 defective is \[P(x=2)=\frac{\binom{3}{2}\binom{9}{2}}{\binom{12}{4}}\]

OpenStudy (anonymous):

clear i hope numerator says number of ways to choose 2 out of 3 times number of way to choose 2 out of 9, denominator is the number of ways to choose 4 out of 12

OpenStudy (anonymous):

ill have a look at it once ive finished writing everything down

OpenStudy (anonymous):

ok armed with this, the rest should be easy because they work the same way

OpenStudy (anonymous):

thank you soooooooooooooooooooooooo much

OpenStudy (anonymous):

will you be around?

OpenStudy (anonymous):

here is another quick example i have a cooler with 6 cans of bud, 4 of miller and 5 of heineken i pick 3 cans at random probability i get 2 buds and a heiny is \[\frac{\binom{6}{2}\binom{5}{1}}{\binom{15}{3}}\]that snappy

OpenStudy (anonymous):

for a while at least

OpenStudy (anonymous):

where did you get 2 out of 9 not defective?

OpenStudy (anonymous):

you are picking 4 total right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if two are defective, how many are not defective?

OpenStudy (anonymous):

the remaining 9

OpenStudy (anonymous):

can you please draw me an image :)

OpenStudy (anonymous):

it would really help

OpenStudy (anonymous):

hold on you are picking 4, and we want the probability that you get 2 defective if you have 2 out of the 4 chosen defective, how many out of the 4 chosen are not defective?

OpenStudy (anonymous):

2 not defective

OpenStudy (anonymous):

right

OpenStudy (anonymous):

and the number of ways to pick 2 defective and 2 not defective in this example is \[\binom{3}{2}\times \binom{9}{2}\]

OpenStudy (anonymous):

ohhhhhhhhhhhh now i get it

OpenStudy (anonymous):

the one on the left is the number of ways to pick two out of three, the one on the right is the number of ways to pick 2 out of 9

OpenStudy (anonymous):

eureka

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

so suppose x=1

OpenStudy (anonymous):

so without any trouble you should be able in a second to write down how to find \[P(x=1)\]

OpenStudy (anonymous):

1 defective, 3 not defective

OpenStudy (anonymous):

right how many ways?

OpenStudy (anonymous):

(3 1) (9 3) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if it was 3 defective, 1 not defective

OpenStudy (anonymous):

so \[P(x=1)=\frac{\binom{3}{1}\binom{9}{3}}{\binom{12}{4}}\]

OpenStudy (anonymous):

(3 3) (9 1)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes including the denominator

OpenStudy (anonymous):

now why the denominator?

OpenStudy (anonymous):

which is trivial to compute since \(\binom{3}{3}=1\) and \(\binom{9}{1}=9\)

OpenStudy (anonymous):

why the denominator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how many light bulbs total?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

how many are you choosing out of that 12?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

so how many ways can you choose 4 out of 12?

OpenStudy (anonymous):

alright so to find the denominator we use binomial coefficient

OpenStudy (anonymous):

???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is therefore \(\binom{12}{4}\) for all of them

OpenStudy (anonymous):

that is why i computed \(\binom{12}{4}\) first

OpenStudy (anonymous):

man this thing was giving me a nightmare until u came along :)

OpenStudy (anonymous):

i have another nightmare coming soon :)

OpenStudy (anonymous):

but now it is more or less easy right?

OpenStudy (anonymous):

you still have to compute the numbers, but i would wolfram for that

OpenStudy (anonymous):

yeah bit after bit im getting to understand the material thanks to you